OpenStudy (anonymous):

int secx dx? what kind of magic should I do? need a step by step solution ^^

OpenStudy (lgbasallote):

\[\large \int \sec x dx\] \[\large \int \sec x \times \frac{\sec x + \tan x}{\sec x + \tan x} dx\]

OpenStudy (mimi_x3):

Write it as: \[\int\limits\frac{secx}{1}dx \] Then multiple top and bottom by \(secx+tanx\) then u-sub.

OpenStudy (lgbasallote):

you will get \[\large \int \frac{(\sec^2 x + \sec x \tan x)dx}{\sec x + \tan x}\] you get it so far right @GuilhermCosta ??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I just don't know what to 'u-sub' =/

OpenStudy (mimi_x3):

let \(u=secx+tanx\)

OpenStudy (lgbasallote):

now if i let u = \(\sec x + \tan x\) du = \((\sec x \tan x + \sec^2 x)dx\) which is our numerator...therefore it becomes \[\large \int \frac{du}{u}\]

OpenStudy (anonymous):

*-* thank you so much!

OpenStudy (anonymous):

we can't mark more than one good answer anymore? o.O

OpenStudy (lgbasallote):

yeah..makes you have to choose medal wisely

OpenStudy (lgbasallote):

apparently some people were abusing the multiple medal system and devaluing the medals

OpenStudy (anonymous):

=/ thank you 2 anyway ^^